# Bridging Information Gaps with Comprehensive Answers: Improving the Diversity and Informativeness of Follow-Up Questions

### v1

@@ Abstract
Effective conversational systems are expected to dynamically generate contextual follow-up questions to elicit new information while maintaining the conversation flow. While humans excel at asking diverse and informative questions by intuitively assessing both obtained and missing information, existing models often fall short of human performance on this task. To mitigate this, we propose a method that generates diverse and informative questions based on targeting unanswered information using a hypothetical LLM-generated “comprehensive answer”. Our method is applied to augment an existing follow-up questions dataset. The experimental results demonstrate that language models fine-tuned on the augmented datasets produce follow-up questions of significantly higher quality and diversity. This promising approach could be effectively adopted to future work to augment information-seeking dialogues for reducing ambiguities and improving the accuracy of LLM answers [^1] Code available at https://github.com/zheliu92/nlp_followupqg_public/.
@@
Zhe Liu* and Taekyu Kang* and Haoyu Wang and Seyed Hossein Alavi and Vered Shwartz University of British Columbia {zheliu92, salavis, vshwartz}@cs.ubc.ca {davidk15, macdude}@student.ubc.ca
@@
{strip}
Figure 1: Illustration of a learner’s cognitive process in generating follow-up questions.
Green: Implicit cognitive stages; Callouts: Explicit (question, answer, follow-up) triplets collected for the dataset.
@@ 1 Introduction
Asking questions is a fundamental mechanism for humans to acquire new information, particularly when existing information is incomplete. While large language models (LLMs) excel at passively answering user queries, their ability to proactively guide conversations, by identifying and addressing gaps in information, remains underdeveloped Liu et al. (2025). Question generation (QG) has gained significant attention in NLP, particularly for its role in improving information-seeking dialogue systems Chen et al. (2024). By generating effective questions, these dialogue systems can make information-seeking more accurate and efficient Qi et al. (2020), resolve ambiguities Li et al. (2017), and ultimately better understand the user’s needs, providing suitable assistance across various domains, including education Laban et al. (2022) and healthcare Arslan et al. (2024); Li et al. (2024).
@@
While most existing QG tasks focus on generating questions directly answerable by a given context Zhao et al. (2018); Pan et al. (2020); Ghanem et al. (2022), which diverges from the human cognitive process of inferring and pursuing missing information, Meng et al. (2023) proposes that models must generate follow-up questions that build on—but are not answerable by—an initial question-answer pair. FollowupQG is more reasonable, but Meng et al. (2023) found that existing models struggle to replicate this human behavior, often producing repetitive or context-bound questions that fail to target unexplored information.
This core challenge can be formulated into two dimensions: (1) identifying information gaps, the unanswered aspects of the initial question, and (2) generating diverse questions that address these gaps. Traditional QG methods Zhao et al. (2018); Pan et al. (2020); Ghanem et al. (2022), focusing on rephrasing or extracting information directly from the input, are less effective at generating questions that infer missing or implied content. In contrast, recent work Mazzaccara et al. (2024); Liu et al. (2025) that attempts to generate follow-up questions that seek missing information lacks explicit mechanisms to model gaps or ensure diversity, resulting in questions that regurgitate existing context or exhibit limited novelty.
To mitigate these limitations, we propose an information gap-driven pipeline that generates follow-up questions by contrasting original answers with synthetically generated comprehensive answers. Our key insight is that comprehensive answers—hypothetical, LLM-generated “complete” responses to the initial question—reveal potential gaps when compared to the original, often incomplete answers. By analyzing these gaps, our method explicitly targets unanswered information, aligning with human cognitive strategies. For example, in Figure 1, if the original answer to “how do clouds form?” explains “clouds form when water vapor cools,” a comprehensive answer might add “…and condenses around dust particles,” exposing the gap “What role do particles play in cloud formation?”.
@@
Specifically, we use GPT-4 (4o: 2024-02-15 -preview) to generate both the comprehensive answers and the information gap-driven follow-up questions and verified the high quality of the generated questions. We then augment the original FollowupQG training set with more than 25,000 synthetic examples (approximately 10$\times$ the original size) and fine-tuned several language models on both the original dataset and our augmented dataset. Instead of using GPT-4 directly for follow-up question generation, we adopt a distillation approach: leveraging GPT-4 to generate high-quality training data, and then fine-tuning smaller models to achieve strong performance at a significantly lower cost. The extensive experimental results demonstrate significant improvements of the augmented dataset over the baselines, both in terms of quality (validity, relevance, informativeness, etc.) and diversity. Our contributions are as follows:
@@
  • We propose a novel approach that generates follow-up questions by identifying information gaps through contrastive analysis of original and generated comprehensive answers.
@@
  • We augment the FollowupQG training set with 25,000 synthetic examples, which achieved 94% validity according to human evaluation.
@@
  • Experimental results show that models fine-tuned on our augmented dataset outperform baselines in diversity and informativeness.
@@ 2 Related Work
Question generation (QG) focuses on automatically generating semantically meaningful and well-structured questions based on a given text Ali et al. (2010). While traditional QG techniques have made significant strides in domains such as machine comprehension Du et al. (2017); Uto et al. (2023), e-commerce Wang et al. (2021), and education Luo et al. (2024), they primarily generate questions based on known answers. This approach contrasts sharply with human questioning behavior, which actively seeks new information from various perspectives. This limitation has led to the emergence of follow-up QG, a task whose goal is to generate questions that explore previously unanswered or underexplored aspects of a given text.
Follow-up QG has evolved from simpler methods, such as template-based and retrieval-driven approaches Kumar and Joshi (2017); Soni and Roberts (2019); B et al. (2020), to more advanced techniques that prioritize informativeness Majumder et al. (2021); Mazzaccara et al. (2024). Knowledge-enhanced approaches, like those in Ge et al. (2023) and Gupta et al. (2022), leverage entity-relation pairs and knowledge graphs to improve the depth of the generated questions. Further advancing this, Liu et al. (2025) combined knowledge graphs with LLMs to increase question informativeness. Efforts to model human-like questioning behavior, such as InquisitiveQG Ko et al. (2020), have relied on crowd-sourced follow-up questions written for news articles rather than those naturally generated by humans, leading to a lack of depth and cognitive diversity.
@@
Meng et al. (2023) found that models often produce questions that are either repetitive or fail to target unexplored information, thus lacking the cognitive diversity and variability seen in human questioning strategies Sultan et al. (2020). While follow-up QG has made significant progress, existing approaches largely focus on generating questions directly, using various model architectures and knowledge enhancement techniques. Our work, however, takes a novel approach inspired by the human cognitive process that directly models information gaps and uses them to guide the follow-up question generation.
@@ 3 Data Augmentation
Effective follow-up question generation (QG) requires models to infer and target gaps between the provided answer and the broader context of a conversation. We follow the task definition presented by the FollowupQG Meng et al. (2023): “to generate follow-up questions that seek new information given the initial question and answer”. For simplicity, we denote the “initial question” as IQ, “initial answer” as IA, and the “follow-up question” as FQ. Critical limitations in the training dataset are identified, including quality issues, which are addressed through dataset cleaning (Sec 3.1). The small scale (2,790 instances) and low diversity of the dataset are tackled by a novel data augmentation pipeline introduced in this paper (Sec 3.2). Specifically, we augment the FollowupQG training set with synthetic data generated by a pipeline that mimics human-like gap-driven questioning. This approach utilizes LLM-generated comprehensive answers to systematically identify missing information in initial answers and generate follow-up questions targeting those gaps. We demonstrate that the augmented data retains high quality (Sec 3.3).
@@ 3.1 Data Cleaning
The FollowupQG dataset is limited by its small scale, comprising 3,790 samples: 2,790 for training, 500 for validation, and 500 for testing. Within the 2,790 training instances, there are only 2,651 unique (IQ, IA, FQ) triplets, indicating duplication. Additionally, the number of 2,648 unique (IQ, IA) pairs suggest minimal follow-up question diversity, as 99.8% of pairs have only one reference FQ. Further analysis also uncovered data quality issues, likely stemming from automated data collection (see Appendix A). To improve the data quality, we did the following:
  • Deduplication. We removed 139 duplicate (IQ, IA, FQ) triplet instances.
  • Reference quality check. We manually filtered out 84 instances where the reference FQ diverged entirely from the initial question.
  • Sensitive content removal. We excluded 24 instances involving topics like self-harm or crime, which LLMs are likely to refuse to answer.
The cleaned dataset (2,543 instances) retained broad topic coverage (2,533 unique question-answer pairs).
@@ 3.2 Augmentation Pipeline
As discussed in Section 3.1, the limited scale of the dataset and the lack of follow-up question diversity hinder the coverage of diverse questioning strategies, restricting model generalization. To address this, we design a GPT-4-based pipeline that augments the original dataset by generating additional follow-up questions. Our pipeline simulates human reasoning through three interconnected stages: comprehensive answer generation, information gap identification, and follow-up question generation [^2] Please refer to Appendix B for the LLM prompts used for the following stages..
@@
To identify gaps in the IA, we generate a comprehensive answer (CA) that represents a complete and thorough response to the IQ. As shown in Figure 2, we prompt GPT-4 iteratively to generate answers to IQ that target different perspectives, such as technical, ethical, and practical, and synthesize a unified CA.
@@
The next step is to identify key concepts or details discussed in the comprehensive answer (CA) but not covered in the initial answer (IA). This is done by prompting GPT-4. For example, in the example shown in Figure 2, the initial answer covers the topic of privacy issues but does not cover areas of cyber security (i.e. an information gap).
@@
Using the identified information gaps, we prompt GPT-4 to generate follow-up questions that address those gaps while maintaining contextual relevance to the IQ and IA. The generated questions must meet three criteria: be (1) answerable by the CA, (2) unanswerable by the IA, and (3) grounded in terminology and context from the IQ.
To match the format of the original follow-up questions in the FollowupQG dataset, we automatically reformat the generated FQs to remove artifacts such as bullets or numbering. Each (IQ, IA) pair is enriched with multiple follow-up questions generated by our pipeline, averaging 10.95 questions per pair. The restructured dataset merges synthetic questions with cleaned human-generated examples, resulting in 27,874 samples—a 10$\times$ increase from the original dataset size. This expanded dataset captures the open-ended nature of human questioning, providing models with diverse and explicit signals to learn strategies for addressing information gaps effectively.
@@ 3.3 Augmented Data Validation
To assess the quality of the generated follow-up questions, we conducted a human evaluation study using Cloud Connect. To ensure high-quality annotations, we restricted participation to native English-speaking annotators with a minimum of 1,000 completed annotation tasks and an approval rating exceeding 90%. A randomly sampled subset of 100 (IQ, IA, FQ) triplets was evaluated based on three key criteria: (1) whether the follow-up question was a valid question [^3] A valid question must be in a question format and ask meaningful information, including Wh-questions (what/why/where/etc.), open-ended questions, probing questions and etc. Meng et al. (2023), (2) whether any component of the triplet contained sensitive information, and (3) the degree of relatedness between the follow-up question and the initial question-answer pair. The full survey format, including example annotations, is provided in Appendix C. The results show that 94% of the follow-up questions are labeled as valid, 92% as not sensitive, and 91% are related to the original (IQ, IA) pair. Inter-annotator agreement was moderate, with a Cohen’s Kappa score of $\kappa=0.73$ McHugh (2012).
@@ 4 Experiment Setup
To evaluate the effectiveness of our proposed pipeline and augmented dataset, we train several variants of the same model on different data. We use BART-large Lewis et al. (2020) as our base model, motivated by its strong performance on the work of FollowupQG Meng et al. (2023). We report the performance of three variations trained on FollowupQG training set. The ORG variant follows the method presented in Meng et al. (2023), trained on the 2,790 original instances from the training set, serving as the baseline model. It employs a seq2seq model that conditions on IQ and IA (IQ <SEP> IA) to generate FQ. [^4] As the implementation from Meng et al. (2023) was not publicly available, we tried to reproduce the performance based on the hyperparameters reported in their paper, however, we had to reduce the learning rate from 5e-5 to 2e-5 avoid training instability. The details about the hyperparameters and difference in performance are available in Appendix D.
We similarly train the FULL variant on all 27,874 instances in the augmented dataset. Finally, we also train the AUG variant on a random sample of 2,790 of the GPT-generated questions (Sec 3.2), to isolate the impact of our data augmentation technique from the effects of increased dataset size. For all variants, we used identical hyperparameters and evaluated them on the original FollowupQG validation and test sets [^5] While other follow-up question generation methods exist, we do not include direct comparisons due to their fundamentally different, non-comparable setups (e.g., Mazzaccara et al. (2024)) and the very recent release of concurrent work (e.g., Liu et al. (2025)), making direct comparison neither straightforward nor feasible. However, we discuss their relevance in Section 2..
@@
To generate diverse but contextually relevant follow-up questions, we input the initial question and answer into the model in the following format: IQ <SEP> IA, and generate 10 follow-up questions by applying beam search with a beam width of 20, selecting the top 10 candidates. We add a diversity penalty of 10 to encourage unique outputs across the groups and set the temperature to $t=1.0$ to maintain a balance between diversity and coherence. The maximum length for each generation is set to 1024 tokens. Duplicate generations are removed.

### v2

@@ Abstract
Generating diverse follow-up questions that uncover missing information remains challenging for conversational agents, particularly when they run on small, locally hosted models. To address this, we develop an information-gap-driven knowledge distillation pipeline [^1] Code available at https://github.com/zheliu92/nlp_followupqg_public/ in which a teacher LLM generates a comprehensive answer, contrasts it with the initial answer to identify information gaps, and formulates gap-bridging follow-up questions. Using this pipeline, we augment the existing FollowupQG dataset tenfold. We then fine-tune smaller student models on the augmented dataset to distill the teacher’s knowledge. Experiments with selected teacher-student model pairs show that fine-tuned students achieve significantly higher informativeness and diversity than variations trained on the original dataset. These findings indicate that our pipeline, which mirrors the human cognitive process of information seeking, provides an efficient distillation channel from state-of-the-art LLMs to smaller models, enabling resource-constrained conversational systems to generate more diverse and informative follow-up questions.
@@
Zhe Liu1*   Taekyu Kang1*   Haoyu Wang1   Seyed Hossein Alavi1,2   Vered Shwartz1,2 1 University of British Columbia   2 Vector Insitute {zheliu92, salavis, vshwartz}@cs.ubc.ca {davidk15, macdude}@student.ubc.ca
@@ 1 Introduction
Asking questions is a fundamental mechanism for humans to acquire new information, particularly when existing information is incomplete. While large language models (LLMs) excel at passively answering questions from users, their ability to proactively guide conversations by identifying and addressing information gaps remains underdeveloped Liu et al. (2025), with smaller models performing even worse. Therefore, the task of question generation (QG) has become a focal point in natural language processing (NLP) for its role in improving information-seeking dialogue systems Chen et al. (2024)—including, making information seeking more accurate and efficient Qi et al. (2020), resolving ambiguities Li et al. (2017), and ultimately better understanding users’ needs to provide suitable assistance across various domains Laban et al. (2022); Arslan et al. (2024); Li et al. (2024).
@@
While most existing QG tasks focus on generating questions directly answerable from a given context Zhao et al. (2018); Pan et al. (2020); Ghanem et al. (2022)—a process that diverges from how humans infer and pursue missing information, Meng et al. (2023) propose the FollowupQG task, which requires models to generate follow-up questions that build on, but are not answerable by, the initial question-answer pair. They create the FollowupQG dataset and show that existing models often produce repetitive or context-bound questions that fail to target unexplored information Meng et al. (2023). The core challenges of the FollowupQG task can be formulated into two dimensions: (1) identifying information gaps, the unanswered aspects of the initial question, and (2) generating diverse questions that target these gaps.
Building on traditional QG methods Zhao et al. (2018); Pan et al. (2020); Ghanem et al. (2022), recent work attempts to generate information-seeking follow-up questions using preference optimization Mazzaccara et al. (2024) and knowledge graphs Liu et al. (2025), but still lack explicit mechanisms to model gaps or ensure diversity. To address these limitations, we propose an information-gap-driven teacher-student knowledge distillation pipeline. In our approach, a teacher LLM generates a hypothetical “complete” response to the initial question, contrasts it with the often incomplete initial answer to identify information gaps, and formulates gap-bridging follow-up questions. By generating multiple follow-up questions, each targeting some unanswered information, this pipeline ensures the diversity and informativeness of the follow-up questions. For example, in Figure 1, if the initial answer to “how do clouds form?” is “clouds form when water vapor cools,” a comprehensive answer might add “… and condenses around dust particles,” which explicitly exposes the gap through contrast and leads to an informative follow-up question such as “What role do particles play in cloud formation?
@@
Our pipeline can be applied across different teacher-student model pairs. In this work, we use GPT-4o (2024-02-15-preview) as the teacher model and BART-large as the student model to verify the pipeline. Specifically, we use GPT-4o to generate the comprehensive answers and follow-up questions. After verifying the quality of the follow-up questions via human evaluation, we then augmented the original FollowupQG training set tenfold and fine-tuned BART-large on both the original dataset and our augmented dataset. Leveraging GPT-4o to generate high-quality training data, and then distilling the teacher’s knowledge into smaller models, our approach achieves strong performance at a significantly lower cost. The experimental results demonstrate significant improvements of the augmented dataset over the baselines, both in terms of quality (validity, relevance, informativeness, etc.) and diversity. Our contributions are as follows:
@@
  • We propose an information-gap-driven teacher-student knowledge distillation pipeline that generates follow-up questions through contrastive analysis of initial answers and generated comprehensive answers.
@@
  • We augment the FollowupQG training set with over 25,000 high-quality synthetic examples.
@@
  • Experimental results show that small models fine-tuned on our augmented dataset outperform peer small-model baselines and achieve near parity with representative LLM-based Teacher and Chain-of-Thought models.
@@ 2 Related Work
Question generation (QG) focuses on automatically generating semantically meaningful and well-structured questions based on a given text Ali et al. (2010). While traditional QG techniques have made significant strides in domains such as machine comprehension Du et al. (2017); Uto et al. (2023), e-commerce Wang et al. (2021), and education Luo et al. (2024), they primarily generate questions based on known answers. This approach contrasts sharply with human questioning behavior, which actively seeks new information from various perspectives. This limitation has led to the emergence of FollowupQG, a task whose goal is to generate questions that explore previously unanswered or underexplored aspects of a given text.
FollowupQG has evolved from simpler methods, such as template-based and retrieval-driven approaches Kumar and Joshi (2017); Soni and Roberts (2019); B et al. (2020), to more advanced techniques that prioritize informativeness Majumder et al. (2021); Mazzaccara et al. (2024). Knowledge-enhanced approaches, like those in Ge et al. (2023) and Gupta et al. (2022), leverage entity-relation pairs and knowledge graphs to improve the depth of the generated questions. Further advancing this, Liu et al. (2025) combined knowledge graphs with LLMs to increase question informativeness. Efforts to model human-like questioning behavior, such as InquisitiveQG Ko et al. (2020), have relied on crowd-sourced follow-up questions written for news articles rather than those naturally generated by humans, leading to a lack of depth and cognitive diversity.
@@
Meng et al. (2023) found that models often produce questions that are either repetitive or fail to target unexplored information, thus lacking the cognitive diversity and variability seen in human questioning strategies Sultan et al. (2020). While follow-up QG has made significant progress, existing approaches largely focus on generating questions directly, using various model architectures and knowledge enhancement techniques Ge et al. (2023); Liu et al. (2025). Our work, however, takes a novel approach inspired by the human cognitive process that models information gaps and uses them to guide the follow-up question generation.
@@ 3 Data Augmentation
Effective FollowupQG requires models to infer and target gaps between the provided answer and the broader context of a conversation. Following the task definition by Meng et al. (2023): “to generate follow-up questions that seek new information given the initial question and answer”, we denote the “initial question” as IQ, “initial answer” as IA, and the “follow-up question” as FQ. We identified critical limitations in the training dataset, including quality issues, which we addressed through dataset cleaning (§3.1). To overcome the small scale (2,790 instances) and low diversity of the dataset, we present a novel data augmentation pipeline (§3.2). Finally, we demonstrate that the augmented dataset retains high quality (§3.3).
@@ 3.1 Data Cleaning
The FollowupQG dataset is limited by its small scale, comprising 3,790 samples: 2,790 for training, 500 for validation, and 500 for testing. Within the 2,790 training instances, there are only 2,651 unique (IQ, IA, FQ) triplets, indicating duplication. Moreover, the dataset consists of 2,648 unique (IQ, IA) pairs, meaning that 99.8% of the (IQ, IA) pairs have only one reference FQ. Training models on this set could thus lead to poor follow-up question diversity. Our further analysis also uncovered data quality issues, likely stemming from automated data collection (see Appendix A). To improve the data quality, we did the following:
  • Deduplication: We removed 139 duplicate (IQ, IA, FQ) triplets.
  • Reference quality check: We manually filtered out 84 instances where the reference FQ diverged entirely from the initial question.
  • Sensitive content removal: We excluded 24 instances involving topics like self-harm or crime, which LLMs are likely to refuse to answer.
The cleaned dataset (2,543 instances) retained broad topic coverage, containing 2,533 unique (IQ, IA) pairs.
@@ 3.2 Augmentation Pipeline
As discussed in §3.1, the limited scale of the dataset and the lack of follow-up question diversity hinder the coverage of diverse questioning strategies, restricting model generalization. To address this, we design a GPT-4o-based pipeline that augments the original dataset by generating additional follow-up questions. Our pipeline simulates human reasoning through three interconnected stages: comprehensive answer generation, information gap identification, and follow-up question generation. [^2] Please refer to Appendix B for the LLM prompts used for the following stages.
@@
To identify gaps in the IA, we generate a comprehensive answer (CA) that represents a complete and thorough response to the IQ. As shown in Figure 2, we prompt GPT-4o iteratively to generate answers to IQ that target different perspectives, and synthesize a unified CA. More specifically, GPT-4o was prompted to generate a combination of new and different answers that do not overlap with the other answers, where each answer focuses on a unique aspect not covered in the other generated answers.
@@
The next step is to identify key concepts or details discussed in the comprehensive answer (CA) but not covered in the initial answer (IA). This is done by prompting GPT-4o. As shown in Figure 2, the initial answer covers the topic of privacy issues but does not cover areas of cyber security (i.e. an information gap). To confirm the validity and reasonableness of the comprehensive answers and identified information gaps, we manually evaluated a random sample of comprehensive answers and ensured that they were accurate and reasonable. Examples of comprehensive answers can be seen in Table 7 and Table 23.
@@
Using the identified information gaps, we prompt GPT-4o to generate follow-up questions that address those gaps while maintaining contextual relevance to the IQ and IA. The generated questions must meet three criteria: be (1) answerable by the CA, (2) unanswerable by the IA, and (3) grounded in terminology and context from the IQ.
[…]
@@ 3.3 Augmented Data Validation
To assess the quality of the generated follow-up questions, we conducted a human evaluation study on Cloud Connect, using Meng et al. (2023)’s survey. To ensure high-quality annotations, we restricted participation to native English-speaking annotators with a minimum of 1,000 completed annotation tasks and an approval rating exceeding 90%. A randomly sampled subset of 100 (IQ, IA, FQ) triplets was evaluated based on three key criteria: (1) whether the FQ was a valid question, [^3] Following Meng et al. (2023), a valid question must be in a question format and ask for meaningful information, including Wh-questions (what/why/where/etc.), open-ended questions, probing questions, etc. (2) whether any component of the triplet contained sensitive information, and (3) the degree of relatedness between the FQ and the (IQ, IA) pair. The full survey format, including example annotations, is provided in Appendix C. The results show that 94% of the FQs are labeled as valid, 92% as not sensitive, and 91% are related to the original (IQ, IA) pair. Inter-annotator agreement was moderate, with a Cohen’s Kappa score of $\kappa=0.73$ Cohen (1960).
@@ 4 Experiment Setup
To assess our proposed pipeline and augmented dataset, we fine-tuned BART-large Lewis et al. (2020) (24 layers, 16 attention heads, hidden size = 1024) on several versions of the FollowupQG data Meng et al. (2023), producing three model variants. BART-large is a seq2seq model that conditions on the concatenated IQ and IA to generate an FQ. We chose it as our base model because of its strong performance reported by Meng et al. (2023). As their implementation is not public, we reproduced their training setup (batch = 8, epochs = 10, Adam Kinga et al. (2015)) and found that the original learning rate of 5e-5 caused instability, so we reduced it to 2e-5; all other hyperparameters remain unchanged. [^4] Full hyperparameter details and reproduction results are provided in Appendix D.
We report three variants trained on the FollowupQG training set as shown in Figure 3: the ORG model is trained on the 2,790 original instances from Meng et al. (2023) and serves as our small-model baseline; the AUG model, trained on a size-matched random sample of 2,790 GPT-generated questions from our augmentation pipeline (Sec. 3.2) to isolate the effect of data quality; and the FULL model, trained on the entire 27,874-instance augmented dataset. All variants share identical hyperparameters and are evaluated on the original FollowupQG validation and test splits.
[…]
@@
To generate diverse but contextually relevant follow-up questions, we input the initial question and answer into the model in the following format: IQ <SEP> IA, and generate 10 follow-up questions by applying beam search with a beam width of 20, selecting the top 10 candidates. We added a diversity penalty of 10 to encourage unique outputs across the groups and set the temperature to $t=1.0$ to maintain a balance between diversity and coherence. The maximum length for each generation is set to 1024 tokens. Duplicate generations are removed.
LLM Baselines.
Beyond the three BART variants, we evaluated two GPT-4o baselines that represent current large-model performance. The Teacher setting executes the full information-gap pipeline from §3.2 on each test instance, returning GPT-4o’s comprehensive answer, the identified gaps, and the resulting follow-up questions. The CoT setting applies the same chain-of-thought prompt but suppresses all intermediate reasoning, outputting only the follow-up questions. Duplicate generations are removed as in the BART variants.